20.0 mL of apple juice is added to some water for a total volume of 120.0 mL of drink solution. Calculate the % apple juice.
Answer by fcabanski(1391)

Concentration of the solution is amount of desired substance divided by total amount. In this case the "desired substance" is apple juice. It's 20 mL. The total is 120 mL.


20/120 = 1/6 = .1666666666 = approx. 16.7%